Find an explicit solution of the given initial-value problem. ✓ 3 ✓ 1 - y2 dx - V1 - x? dy = 0, y(O) = 2
Answer #1

Given initial value problem, vi-yz dx-V1-x² dy = 0, y co) = 13 as follows We write, o can Vi-ya dx = V1-x² dy da dy -x² VI-y²sintx = siny tc Taking sin both sides, we get To x = y + sinc find Apply yco)=13 in equation e 2 we get o = 3+ sinc = c = sin